OpenStudy (anonymous):

what is a three step equation in which the answer is 2?

OpenStudy (anonymous):

2x + 3 = 7 Here is a two step equation, can you add the third step?

OpenStudy (anonymous):

4=2+x

OpenStudy (anonymous):

you already asked same question like 6 times, can't you do it yourself?

OpenStudy (mertsj):

Let's start with the number 2 and call it x. Let's multiply it by some number. I choose 6 so now we have 6x. Let's add something to it, maybe 4. So now we have 6x+4. Now we must do a third thing to it so let's divide it by 4. So now we have (6x+4)/4. But we know the secret and the secret is that x is 2. So if we multiply 2 by 6 we get 12. Now add 4. That is 16. Now divide by 4 and the answer is 4. So our equation is (6x+4)/4 =4

OpenStudy (anonymous):

happy with that answer ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

still there?

OpenStudy (anonymous):

yes